# A Unified Formal Theory on the Logical Limits of Symbol Grounding

### v2

@@ Abstract
This paper synthesizes a series of formal proofs to construct a unified theory on the logical limits of the Symbol Grounding Problem. We demonstrate through a four-stage argument that meaning within a formal system must arise from a process that is external, dynamic, and non-algorithmic. First, we prove that any purely symbolic system, devoid of external connections, cannot internally establish a consistent foundation for meaning due to self-referential paradoxes. Second, we extend this limitation to systems with any finite, static set of pre-established meanings, proving they are inherently incomplete. Third, we demonstrate that the very "act" of connecting an internal symbol to an external meaning cannot be a product of logical inference within the system but must be an axiomatic, meta-level update. Finally, we prove that any attempt to automate this update process using a fixed, external "judgment" algorithm will inevitably construct a larger, yet equally incomplete, symbolic system. Together, these conclusions formally establish that the grounding of meaning is a necessarily open-ended, non-algorithmic process, revealing a fundamental, Gödel-style limitation for any self-contained intelligent system.
@@ 1 Introduction
Third, in Section 4, we investigate the nature of this dynamic process. What is the "grounding act" that expands the system’s connection to the world? We prove that this act is non-inferable. No internal, condition-triggered logical rule can deduce a command to ground a new symbol without risking a contradiction where the system’s inference invalidates its own proven theorems. This suggests the grounding act is an axiomatic, meta-level update, not a logical deduction.
@@ 2 The Formal Framework and the Impossibility of Self-Grounding
Any purely symbolic system that satisfies Axioms 2.5 and 2.6 cannot internally provide a complete and consistent determination of groundability for all its propositions.
@@
    By the Diagonal Lemma (Axiom 2.6), there exists a sentence $P_{\text{ungroundable}}$ such that the system proves:
@@ 3 The Incompleteness of Statically Grounded Systems
    Since $G$ is finite, the property of "being groundable by GG" is recursive. Therefore, the system can express a specific predicate $Groundable_{G}(x)$, which is true if and only if the formula with Gödel number $x$ is groundable by the set $G$.
@@
4 The Non-Inferable Nature of the Grounding Act
If the grounding set must expand, what is the nature of this expansion? Is the "grounding act"—the connection of a new symbol to an external experience—a logical deduction? We will now prove that it is not.
@@ 4 The Non-Inferable Nature of the Grounding Act
A Grounding Act $A(s,g)$ is a meta-level operation that updates the system’s definition function, adding a new grounded symbol $g$ to the definitional set of a previously ungrounded symbol $s$.
Theorem 4.1 (The Non-Inferability of the Grounding Act).
In a consistent deductive symbolic system $\mathcal{S}$, there exists no general, internal, and condition-triggered logical rule that can deduce a command to ground any of its provable but ungrounded sentences.
@@
    From Theorem 3.2, we know we can construct a sentence $P_{new}$ such that (a) $\mathcal{S}\vdash P_{new}$ (it is provable) and (b) $P_{new}$ is ungrounded by the current set $G$.
@@
    Assume for contradiction that there exists an internal "Grounding Principle" in $\mathcal{S}$. This principle would state: "For any provable sentence $\phi$, if $\phi$ is ungrounded, then the system must deduce a command to execute the grounding act $A(\phi,g)$."
@@
    Applying this assumed principle to $P_{new}$, the system would deduce a command to ground $P_{new}$. Let’s say it executes $A(P_{new},g)$.
  4. 4.
    This leads to a direct contradiction. After the grounding act, the system is in a new state $\mathcal{S}^{\prime}$.
    • In $\mathcal{S}^{\prime}$, the sentence $P_{new}$ has, by definition, become groundable.
    • However, the content of the sentence $P_{new}$ is "This sentence is not groundable by the set G."
    • This assertion, which was true in $\mathcal{S}$, has now become false in $\mathcal{S}^{\prime}$.
  5. 5.
    The contradiction is as follows: The system $\mathcal{S}$ used its rules of inference to prove a theorem ($P_{new}$) and then used another rule of inference (our assumed Grounding Principle) to deduce an action that directly falsified its own theorem. A consistent logical system cannot have rules that invalidate its own conclusions.
  6. 6.
    Therefore, the initial assumption must be false. The grounding act cannot be the result of logical inference within the system.
This proves that grounding is not an inference but an update. It is not a conclusion reached within the system’s existing rules but is analogous to adding a new axiom from outside. This provides a formal basis for the "Aha!" moments or paradigm shifts in cognition, which are not deductions but restructurings of the axiomatic foundation itself.
@@ 6 Conclusion
    The act of expanding this connection—the grounding act—is not a logical deduction but a non-inferable, axiomatic update to the system’s foundation.
@@ Appendix A Formal Construction of the Groundability Predicate
This appendix provides a more detailed formal construction of the $Groundable(n)$ predicate within a system containing Peano Arithmetic, demonstrating that Axiom 2.5 is well-founded.

### v3

@@ Abstract
This paper synthesizes a series of formal proofs to construct a unified theory on the logical limits of the Symbol Grounding Problem. We demonstrate through a four-stage argument that meaning within a formal system must arise from a process that is external, dynamic, and non-algorithmic. First, we prove that any purely symbolic system, devoid of external connections, cannot internally establish a consistent foundation for meaning due to self-referential paradoxes. Second, we extend this limitation to systems with any finite, static set of pre-established meanings, proving they are inherently incomplete. Third, we demonstrate that the grounding process is logically incomplete; specifically, the ’act’ of connecting internal symbols to novel, emergent external meanings cannot be a product of logical inference within the system but must be an axiomatic, meta-level update. Finally, we prove that any attempt to automate this update process using a fixed, external "judgment" algorithm will inevitably construct a larger, yet equally incomplete, symbolic system. Together, these conclusions formally establish that the grounding of meaning is a necessarily open-ended, non-algorithmic process, revealing a fundamental, Gödel-style limitation for any self-contained intelligent system.
@@ 1 Introduction
Third, in Section 4, we investigate the nature of this dynamic process. We prove that the grounding process as a whole is logically incomplete. While some routine connections might be automated, we prove that the crucial act of grounding novel, system-transcending truths is non-inferable. No general internal rule can govern this process without risking contradiction, suggesting that these key grounding acts are axiomatic, meta-level updates, not logical deductions.
@@ 2 The Formal Framework and the Impossibility of Self-Grounding
Any purely symbolic system that satisfies Axioms 2.1 and 2.2 cannot internally provide a complete and consistent determination of groundability for all its propositions.
@@
    By the Diagonal Lemma (Axiom 2.2), there exists a sentence $P_{\text{ungroundable}}$ such that the system proves:
@@ 3 The Incompleteness of Statically Grounded Systems
    Since $G$ is finite, the property of "being groundable by GG" is recursive. Therefore, the system can express a specific predicate $Groundable_{G}(x)$, which is true if and only if the formula with Gödel number $x$ is groundable by the set $G$. (See Appendix A).
@@
4 The Logical Incompleteness of the Grounding Act
If the grounding set must expand, what is the nature of this expansion? While some "routine" grounding acts might be automated by internal logic, is the grounding process as a whole reducible to logical deduction? We will now prove that it is not, by showing that there exists no general logical rule that can handle all grounding acts, and therefore, non-inferable grounding acts must exist.
@@ 4 The Logical Incompleteness of the Grounding Act
A Grounding Act $A(s,g)$ is a meta-level operation that updates the system’s definition function, connecting an internal symbol $s$ to a grounded symbol $g$. This act can take two forms:
  1. 1.
    G-Expansion: $g$ is a new symbol not previously in $S$ or $G$, representing a new external experience. $G$ is updated to $G^{\prime}=G\cup\{g\}$.
  2. 2.
    G-Internal Connection: $g$ is an existing symbol already in $G$. The definition function $D$ is updated, e.g., $D(s)^{\prime}=D(s)\cup\{g\}$.
Theorem 4.1 (Incompleteness of Internal Logical Grounding).
In a consistent deductive symbolic system $\mathcal{S}$, there exists no general, internal, and condition-triggered logical rule ($\vdash$) that can deduce a command to execute a grounding act for all provable but ungrounded sentences.
@@
We must demonstrate that no such general internal rule can exist, by analyzing both forms of the grounding act.
@@
    Case 1: Grounding via G-Expansion ($g\notin G$).
    Assume for contradiction that a general internal rule $\vdash$ can deduce the command $A(s,g)$, where $g$ is a new symbol representing an external experience not yet in $S$ or $G$.
    This is an information-theoretic contradiction. Logical inference ($\vdash$) is a closed computational process that operates only on the axioms and symbols already defined within the system $\mathcal{S}=\langle S,D,G,\vdash\rangle$.
    The symbol $g$, representing new external information, is by definition outside of this formal universe. The system’s internal rules cannot deduce a conclusion that contains a symbol ($g$) which it cannot access, name, or know. Therefore, any act that expands $G$ must be an external, axiomatic update, not an internal inference.
@@
    Case 2: Grounding via G-Internal Connection ($g\in G$).
    This is the more subtle case. Assume for contradiction that there exists a general, internal "Grounding Principle" rule in $\mathcal{S}$. This rule would state:
    "For any provable sentence $\phi$, if $\neg Groundable_{G}(\ulcorner\phi\urcorner)$ is true (i.e., $\phi$ is ungrounded), then the system must deduce ($\vdash$) a command to execute $A(\phi,g)$ for some relevant, existing $g\in G$."
    Now, let us apply this assumed general rule to the sentence $P_{new}$ from Theorem 3.1.
    • Initial State ($\mathcal{S}$): The system has proven $\mathcal{S}\vdash P_{new}$. The premise of the rule $\neg Groundable_{G}(\ulcorner P_{new}\urcorner)$ is true.
    • Action: The (assumed general) Grounding Principle rule fires. The system deduces a command and executes the grounding act $A(P_{new},g_{existing})$, where $g_{existing}$ is some symbol already in $G$. This updates the definition function $D$ such that $g_{existing}\in C(P_{new})$.
    • New State ($\mathcal{S}^{\prime}$): The system is now in state $\mathcal{S}^{\prime}$. Crucially, the grounding set itself is still G.
    • The Contradiction: We re-evaluate the predicate $Groundable_{G}(\ulcorner P_{new}\urcorner)$ in state $\mathcal{S}^{\prime}$.
      • The definitional closure $C(P_{new})$ now contains $g_{existing}$.
      • The intersection $C(P_{new})\cap G$ is no longer empty, because $g_{existing}$ is in both sets.
      • Therefore, the predicate $Groundable_{G}(\ulcorner P_{new}\urcorner)$ is now true.
    • The content of the sentence $P_{new}$ (which is $\neg Groundable_{G}(\ulcorner P_{new}\urcorner)$) has thus become false.
@@
    Conclusion of Proof: We have a formal contradiction. The system $\mathcal{S}$ used an assumed general internal rule of inference ($\vdash$) to deduce an action that directly falsified one of its own proven theorems ($P_{new}$).
    A consistent logical system cannot contain an internal rule that invalidates its own conclusions. This form of self-referential paradox is structurally analogous to the proof of Gödel’s Second Incompleteness Theorem.
    Both potential forms of a general grounding rule are thus shown to be impossible. Case 1 fails on information-theoretic grounds, and Case 2 (the general internal rule) fails on grounds of logical inconsistency.
    This does not mean that no grounding act can be deduced (a restricted, specific rule for a limited subset of symbols may well be possible). However, it proves that the grounding process as a whole is not reducible to internal logic. Because no general rule exists that can handle all cases, it follows logically that there must exist grounding acts (like for $P_{new}$) that are non-inferable from within the system.
This proves that while some grounding acts may be deductive, the grounding process in its entirety is not. The grounding of novel, system-transcending truths (like $P_{new}$) is not a conclusion reached within the system’s existing rules but is analogous to adding a new axiom from outside. This provides a formal basis for the "Aha!" moments or paradigm shifts in cognition, which are not deductions but restructurings of the axiomatic foundation itself.
@@ 6 Conclusion
    The act of expanding this connection for novel, emergent truths is not a logical deduction. This proves that the grounding process is logically incomplete, and that non-inferable, axiomatic updates are a necessary part of it.
@@ Appendix A Formal Construction of the Groundability Predicate
This appendix provides a more detailed formal construction of the $Groundable(n)$ predicate within a system containing Peano Arithmetic, demonstrating that Axiom 2.1 is well-founded.